### v3

# Do Text-to-Vis Benchmarks Test Real Use of Visualisations?

Abstract
Large language models are able to generate code for visualisations in response to user requests. This is a useful application, and an appealing one for NLP research because plots of data provide grounding for language. However, there are relatively few benchmarks, and it is unknown whether those that exist are representative of what people do in practice. This paper aims to answer that question through an empirical study comparing benchmark datasets and code from public repositories. Our findings reveal a substantial gap in datasets, with evaluations not testing the same distribution of chart types, attributes, and the number of actions. The only representative dataset requires modification to become an end-to-end and practical benchmark. This shows that new, more benchmarks are needed to support the development of systems that truly address users’ visualisation needs. These observations will guide future data creation, highlighting which features hold genuine significance for users.
1 Introduction
Text-to-Vis is the task of receiving a request for a visualisation expressed in human language and some data as input and generating code that will produce the visualisation. A system with this ability would enable faster and more complex data analysis, but there are relatively few benchmark datasets. Those that do exist either focus on generating a single response Luo et al. (2021); Srinivasan et al. (2021); Chen et al. (2021), or consider dialogue, but with limited flexibility in code Shao and Nakashole (2020); Song et al. (2023). Most of these datasets used generated data. The space of code variation was defined by researchers, and researchers were only involved in verifying the data. This raises the question of whether these datasets are representative of real-world use of data visualisations.
In this study, we gathered publicly available code from the Stack [^1] A 6TB collection of open source code from GitHub Kocetkov et al. (2022) to analyse human preferences in making visualisations using libraries across four programming languages: Python, R, Javascript, and Vega. Since each library has different names for the same visualisation types and properties, we extracted key visualisation code and developed a cross-language mapping for several hundred functions and arguments.222 For example, a bar plot is produced with bar() and barh() in Python, but barplot() in R. Our data and code are at https://github.com/giahy2507/text-to-vis-bench-assessment.
Using this aligned data, we analysed user behaviours when making visualisation and identified similarities and differences between real-world and benchmark datasets. Our analysis considered the chart types, functions called to define properties, and the arguments that modify how those functions behave. We observed that (1) existing benchmarks tend to focus on only one aspect of the text-to-vis challenge, either code synthesis, data presentation, or aesthetic attribute adjustment, and (2) only one of the datasets is consistent with real-world use, and that dataset is limited by its lack of executability of code outputs.
Success on existing benchmarks does not mean systems are useful for real-world use. We need new benchmarks that cover all aspects of the problem and are consistent with the patterns of use. Only then will we be able to measure our progress on this valuable and challenging task.
2 Related Work
Recently, a common approach to creating text-to-visualisation datasets involves the automatic synthesis of visualisations followed by human intervention for annotation Luo et al. (2021); Shao and Nakashole (2020); Srinivasan et al. (2021); Song et al. (2023); Chen et al. (2021). Although this approach is straightforward, datasets produced using this method often contain inherent problems. For instance, nvBench, the largest benchmark dataset for this task, was synthesized from Spider Yu et al. (2018), a text-to-SQL dataset containing several limitations Suhr et al. (2020), and was only partially reviewed by novices and experts for quality assurance, resulting in numerous issues Li et al. (2024). Similarly, ChartDialog contains limitations as the data for visualisation were automatically generated. These shortcomings underscore the need for improved methodologies in dataset creation to ensure their validity and usability.
Recent research has prioritized ecological validity to enhance benchmark dataset quality across various domains, aiming to align them with real-world applications De Vries et al. (2020); Qi et al. (2023); Lu et al. (2023). Ensuring that the data used to train and test models accurately reflects users’ objectives in practical scenarios is crucial. However, achieving ecological validity remains challenging in text-to-vis. Despite efforts in benchmark creation, such as involving humans in the process for nvBench and ChartDialog, their effectiveness in representing real-world visualisation tasks and user objectives is still being explored.
3 Data Collection
Instead of examining user preferences through chart images or relying on experts to comprehend how visualisations are made, our approach involves the analysis of publicly available programming codes specifically designed for creating visualisations such as line, bar, and scatter charts. It means we can have a wide range of samples from different programmers and their preferences when making visualisations. We believe that the analysis can provide a broad understanding of the essential components that are widely used.
We used code files from The Stack  [^3] https://huggingface.co/datasets/bigcode/the-stack-dedup to conduct our investigation on four diverse and widely used visualisation libraries. They include Matplotlib [^4] https://matplotlib.org/, Graphics [^5] https://www.rdocumentation.org/packages/graphics, ChartJS [^6] https://www.chartjs.org/docs/latest/, and Vega-Lite [^7] https://vega.github.io/vega-lite/, which were built on top of Python, R, Javascript, and JSON schema, respectively. After downloading, we only selected files containing the initial codes indicative of visualisation library usage (e.g., import matplotlib.pyplot as plt). Finally, we used abstract syntax tree (AST) parsers and heuristics to accurately extract library-related variables, function names, arguments, and explicit values. The details are described in Appendix A, while Table 1 (upper) presents their statistics.
# samples Proportion (%) # functions
nb-Matplotlib 385,338 35.89 6,443,220
py-Matplotlib 464,463 3.5 4,484,368
Graphics 6,721 17.15 53,325
ChartJS 2,714 0.0128 8,847
Vega-Lite 1,093 0.0013 15,664
nvBench 7,241 40,478
ChartDialog 3,284 14,690
PlotCoder 97,706 254,251
Table 1: Statistics of real-world and benchmark data. “Proportion” indicates the proportion of the library’s code in the investigated programming languages. “nb-Matplotlib” indicates code from Jupyter notebooks, while “py-Matplotlib” indicates code from Python files.
As for benchmark datasets, we examined three publicly available ones, including nvBench Luo et al. (2021), ChartDialog Shao and Nakashole (2020), and PlotCoder Chen et al. (2021). They vary in settings and scales, as described in Table 1 and 2. While nvBench and ChartDialog serve as end-to-end text-to-vis benchmarks, involving prompts and plotted data as input as well as visualisation as output, PlotCoder is purely a code synthesis dataset, lacking two of these essential components. Appendix B depicts their examples.
Benchmark Input Output Annotation note
nvBench prompt, data code, vis
Auto-generated based on a
text-to-SQL benchmark
ChartDialog prompt, data, vis code, vis Manually annotated
PlotCoder prompt, code code Auto-extracted
Table 2: Description of benchmark datasets.
4 Cross-language Mapping Table
To compare the data described in previous sections, we established a cross-language mapping table, based on frequently used parameters. This involved selecting the top 500 frequently used parameters, identifying categories and attributes, and finally, clarifying correctness based on the libraries’ documentation and code execution. Ultimately, the mapping table comprises 8 categories, 62 attributes, and around 850 parameters across 4 visualisation languages. Figure 1 briefly shows an attribute named “x-axis title,” while details can be found in Appendix C.
Figure 1: A sample in the cross-language mapping
5 Analysis & Discussion
5.1 Comparison of chart types
Figure 2 (upper) depicts the distribution of four common plot types across real-world datasets and nvBench  [^8] We categorize histograms as bar charts, while polar pie and doughnut charts are grouped as pie charts. Each dataset shows distinct preferences for specific plot types. The distribution of nvBench, a benchmark based on Vega-Lite grammar, is significantly misaligned with that of Vega-Lite, where the bar chart dominates other types, accounting for over 80%, while the remaining are around 7%.
Figure 2 (lower) depicts the distribution of seven plot types across four Python-based datasets. Generally, the distribution between Matplotlib and PlotCoder shows notable similarity. This trend is attributable to the source for data collection, GitHub. In contrast, ChartDialog showcases a varied array of plot types; however, its distribution differs from that of Matplotlib. Specifically, ChartDialog witnesses a scarcity of scatter plots and an overabundance of pie charts, contours, and stream plots.
These findings imply that nvBench and ChartDialog are not testing the same distribution with real-world data in terms of plot types. As suggestions for future dataset makers, it is crucial to tailor the distribution of chart types according to the specific needs and domains of the intended users. Moreover, given the imbalance in both real-world data and benchmarks, a critical consideration is to conduct separate evaluations focusing specifically on plot types, acknowledging their distinct roles.
Figure 2: Plot type distribution over 8 datasets.
5.2 Comparison of attributes
Using the cross-language mapping table and parsed data (function names and arguments), we computed the normalized frequency for 62 attributes within each dataset, as shown in Figure 8 located in the appendix section. We used these frequencies to determine the Spearman’s rank correlation coefficient across eight datasets, as illustrated in Figure 3.
Figure 3: Spearman’s rank correlation coefficient in terms of frequent attributes
The real-world datasets have a significant correlation, with Spearman’s values surpassing 0.7, except for ChartJS, which displays a moderate correlation with coefficients hovering around 0.5. As for the benchmarks, ChartDialog and nvBench show a weak correlation with their direct counterparts, Matplotlib and Vega-Lite, respectively. This means many attributes that were frequently used by end users have not been tested in these benchmarks. These include titles, axes-scale limits, tick labels, opacity, histogram bins, legend visibility, and multiple plots handling, as visualized in Figure 8.
Conversely, PlotCoder demonstrates a strong alignment with real-world data, with Spearman’s values ranging from 0.7 to 0.9. The correlation highlights PlotCoder’s potential as a resource for crafting end-to-end text-to-vis benchmarks. This extension is necessary because PlotCoder lacks plotted data in the input as well as visual output. Moreover, it is an automatically extracted dataset focusing on visualization code. This raises questions about the executability of the code and whether the visual output aligns with the user’s intention after execution.
5.3 Comparison of attributes when permitted
Some attributes can only be activated (or permitted) if specific preconditions are met. For instance, the "bar thickness" attribute can be set if the user plots data on a bar chart and adjusts the width parameter. Consequently, these attributes may appear infrequently in the dataset, but users often specify their preferred values. Therefore, analyzing these attributes is crucial for a deeper understanding of end users’ preferences.
In this analysis, we computed the frequency of attributes for a given visualization type or action (e.g. plt.bar()). This calculation is applied to each attribute in the mapping table and visualized as a heat map in Figure 9 located in Appendix D. We focus solely on examining this behaviour in Python-based datasets, including Matplotlib-nb, Matplotlib-py, PlotCoder, and ChartDialog. This is because nvBench does not prioritize user intention for modifying aesthetic attributes while others have different characteristics.
The Spearman’s coefficient calculation among these datasets reinforces our previous findings in Section 5.2. Matplotlib-nb, Matplotlib-py, and PlotCoder show significant correlations with Spearman’s scores above 0.8, whereas ChartDialog lacks correlation with scores below 0.1. While attributes such as axes’ scales, edge color, marker size, pie chart characteristics, legend labels, and grid line attributes receive considerable attention in ChartDialog, end users less frequently specify them and often rely on the library’s automatic assignment. Additionally, dataset creators should consider attributes such as histogram bins, pie precision digits, error-bar visibility, and annotation attributes, which are highly demanded by end users.
5.4 Comparison of program complexity?
To answer this question, we calculate the average count of distinct visualisation functions and parameters within each code file and present the findings in Table 3. In this section, we omit ChartDialog because it is a slot-filling dataset, which can not be counted for the number of functions or parameters.
Regarding benchmarks, they differ significantly from their direct counterparts by employing much fewer functions and parameters, around half of the numbers. Across most real-world data, users typically employ a sufficient number of functions and parameters, ranging from 3 to 7 functions and 10 to 14 parameters. The top 7 functions used in Matplotlib-py exemplify this, aiming to plot the data, save figures, assign titles, and adjust legends. The higher figures in the Vega-Lite dataset can be justified by its nature as a visualisation language (not a library built on top of a programming language). These findings can be attributed to the development of benchmark datasets, considering the complexity of the visualisation program.
No. Funcs No. Params
py-Matplotlib 6.40 10.61
nb-Matplotlib 6.19 10.54
PlotCoder 4.05 6.03
Graphics 3.20 13.82
ChartJS 6.51 12.08
Vega-Lite 10.73 19.27
nvBench 4.59 10.02
Table 3: Average number of functions and parameters
6 Conclusion
In this paper, we analyzed whether text-to-vis benchmarks accurately reflect real-world usage by presenting analyses of chart types, frequent attributes, and program complexity. Our results show that only one of the standard three benchmarks is aligned with real-world use. That dataset has its own critical limitation: it cannot be used as an end-to-end benchmark, going from a request and data as input to a visualization as output. As well as critiquing current benchmarks, we provide guidance for future benchmark development, suggesting the evaluation of relevant attributes and challenging charts that better reflect end users’ preferences. Such a benchmark would guide the development of useful and impactful systems.
Limitations
This study offers analyses of datasets and acknowledges several limitations. Firstly, our examination was restricted to only four visualisation libraries, each corresponding to a different programming language. This narrow scope may not adequately capture the diversity of applications and use cases within the field. Although we attempted to analyse MatLab code files in The Stack dataset, they are miscategorised in The Stack, processed with the wrong extension. [^9] https://huggingface.co/datasets/bigcode/the-stack-dedup/blob/main/programming-languages.json Despite our efforts to clarify this issue by reaching out to the project authors, we have yet to receive a response. Secondly, our investigation is based on public code, mainly representing programmers with different visualisation levels, including novices, practitioners, and experts. If the target users in a visualisation application like Tableau  [^10] https://www.tableau.com/, the level of validity of our research will decrease. Lastly, this study concludes with an analysis and assessment of existing benchmark datasets without proposing solutions. Nevertheless, we believe that the insights and recommendations provided in this work are valuable for any dataset maker and future studies.
Ethics Statement
The data used in this research can be found publicly in the repositories of the cited papers, GitHub, or HuggingFace. Those who want to use the processed data in our repository will need to follow the terms and conditions of The Stack dataset [^11] https://huggingface.co/datasets/bigcode/the-stack-dedup.
Acknowledgments
This material is partially supported by the Australian Research Council through a Discovery Early Career Researcher Award and the Commonwealth Scientific and Industrial Research Organisation (CSIRO). We extend our gratitude to the anonymous reviewers for their constructive feedback and valuable advice on our submissions.
Appendix A Code Parsing
After obtaining code files for Python and R, we used abstract syntax tree (AST) parsers and heuristics to accurately extract variables, function names, arguments, and explicit values. Subsequently, we tracked the assigned variables to correctly select the functions used in Matplotlib while a list of Graphics’ functions was used to filter for this library.
To extract ChartJS specifications, we initially used an AST parser to extract all JSON data from the Javascript code files. Subsequently, a heuristic selection method was applied to filter JSON containing the three essential components of this library, namely "type," "data," and "options." This is because ChartJS relies on the JSON format as its foundation, serving as the input for executing functions in Javascript.
Vega-Lite can appear in both JSON and Javascript files, as it is a JSON schema visualisation language. Therefore, we utilized the above methods for extraction. In detail, after extracting JSON data from code files, we exclusively extracted snippets containing Vega-Lite schema   [^12] Vega-Lite schema: v1, v2, v3, v4, v5,, which is a mandatory field of vega-lite specification.
After achieving functions, arguments, assigned values, and JSON specifications, targeting the visualisation libraries, we transformed them into a universal format to facilitate more accessible analysis and further processing. For instance, a command in Python ax.plot(x, color=’green’, marker=’o’), which plots a line graph of ‘x’, with marker ‘o’ and colour ‘green’, can be parsed into a JSON as {"func_name": "plot", args: ["x"], kargs: {"color": "green", "marker": "o"}}. An example of translating JSON to universal format can be seen in Figure 4.
Figure 4: The process of converting JSON to universal format
Regarding nvBench and PlotCoder, they contain visualisation code in Vega-Lite and Python, so the process was the same as described above. When it comes to ChartDialog, a slot-filling dataset, we converted each user’s intent to a function with changed slots as keyword parameters. For example, a user’s intent "smaller radius, increase text size" modifying a pie chart is transformed as universal JSON format {’func_name’:’pie’, kargs:{’radius: ’small’, ’font_size’:’large’}}.
Appendix B Benchmarks’ Examples
Figures 56, and 7 illustrate examples for nbBench, ChartDialog, and PlotCoder, respectively.
Figure 5: A sample in nvBench dataset
Figure 6: A sample in ChartDialog dataset
Figure 7: A sample in PlotCoder dataset
Appendix C Cross-language Mapping Table
The procedure for making the cross-language table is as follows. Initially, we compiled the top 100 frequent parameters from the real-world dataset in 4 languages: Matplotlib, Graphics, ChartJS, and Vega-Lite. Subsequently, the parameters were identified into different categories and attributes in a mapping table. The table was further expanded by investigating relevant parameters within the top 500. If a specific language lacked relevant parameters for a given attribute in the top 500 (resulting in a blank cell), we persistently searched through the remaining list until a match was found. Cells where no relevant parameter was identified led to the annotation of “not found.” This identification and verification process includes understanding plotting parameters, identifying them in API documents, asking ChatGPT 131313 https://chat.openai.com https://bard.google.com/chat for explanations and relevant parameters, and executing example codes.
Due to the length, we only include a part of the table in Table 4. The whole one can be found in our repository at https://github.com/giahy2507/text-to-vis-bench-assessment.
Category Attribute
Axes
x-title, y-title, x-y-title-fontsize, x-y-title-color,
x-y-lim, x-y-ticks-labels, x-y-ticks-labels-color,
x-y-ticks-labels-rotation, x-y-scale,
x-y-ticks-fontsize, x-axis-ticks-visible,
y-axis-ticks-visible, x-y-scale-position,
invert-x-y-axis
Data Appearance
filled-color, edge-color, opacity, linewidth,
markersize, linestyle, line-capstyle, markerstyle,
bar-thickness, bar-data-stacking, hist-bins,
pie-explode, pie-label-distance,
pie-percentage-distance, pie-precision-digits,
pie-radius, errbar-cap-size, errbar-cap-thick,
errbar-color, errbar-visible
Annotation
ann-text/label, ann-fontsize, ann-possition,
ann-font
Main title
title, title-fontsize, title-color, title-position,
subtitle, subtitle-fontsize
Legend
legend-title, legend-fontsize, legend-position,
legend-labels, legend-labels-color,
legend-is-display
Grid
grid-visible, grid-color, grid-linestyle,
grid-linewidth
Format size, dpi, saving-format
Other
bounding-box/border, background,
margin/padding, multiple-plots
Table 4: Categories and Attributes in cross-language mapping table
Appendix D Calculation for heat map figures
Figure 8 shows heat maps of the most common visualization attributes over 7 datasets, where the more intense green colour indicates a higher percentage of usage within the dataset. The calculation for each attribute is $k/n$, where:
  • $k$ is the number of times that attribute’s arguments are specified
  • $n$ is the number of times that all arguments are specified
As for the heat map in Figure 9, there are two cases influencing different levels. For attributes impacting the program level, such as title, x-axis title, and x-y tick labels, the percentage is derived from how frequently a program includes arguments for a specific attribute. Conversely, for local attributes affecting the function level, like filled colour, opacity, and bar thickness, the percentage is calculated based on the frequency of functions containing arguments for the given attribute. The calculation is as follows. Let
  • $k$ is the number of times that attribute’s arguments are specified. $p$ is the number of times that attribute’s functions are used
  • $z$ represents the total number of programs in the dataset, while $g$ denotes the number of programs in which the attribute is utilized (any of the attribute arguments is used).
While a figure for a given program-level attribute is $g/z$, that for function-level one is $k/p$.
multiple-plotsmargin-paddingbackgroundbounding-box-bordersaving-formatdpisizegrid-linewidthgrid-linestylegrid-colorgrid-visiblelegend-is-displaylegend-labels-colorlegend-labelslegend-positionlegend-fontsizelegend-titlesubtitle-fontsizesubtitletitle-positiontitle-colortitle-fontsizetitleann-fontann-possitionann-fontsizeann-text-labelerrbar-visibleerrbar-colorerrbar-cap-thickerrbar-cap-sizepie-radiuspie-precision-digitspie-percentage-distancepie-label-distancepie-explodehist-binsbar-data-stackingbar-thicknessmarkerstyleline-capstylelinestylemarkersizelinewidthopacityedge-colorfilled-colorinvert-x-y-axisx-y-scale-positiony-axis-ticks-visiblex-axis-ticks-visiblex-y-ticks-fontsizex-y-scalex-y-ticks-labels-rotationx-y-ticks-labels-colorx-y-ticks-labelsx-y-limx-y-title-colorx-y-title-fontsizey-titlex-titleAttribute2.83.00.10.10.20.98.30.10.20.10.75.00.07.92.30.50.10.20.40.10.01.28.50.10.90.70.50.40.00.00.10.00.10.00.00.11.00.00.43.50.02.91.72.52.40.611.20.10.20.60.90.80.30.70.13.02.30.12.68.38.2Matplotlib-nb2.65.10.10.30.51.65.10.10.20.20.84.50.07.32.40.50.10.20.70.10.00.86.70.21.61.00.50.50.10.00.10.00.10.00.00.00.60.10.53.60.03.31.93.22.40.711.70.10.31.01.31.00.40.50.13.43.50.22.27.26.9Matplotlib-py2.41.70.00.20.10.37.20.00.10.10.74.40.05.81.90.20.00.20.50.00.01.111.50.00.30.20.20.10.00.00.00.00.30.00.00.12.70.00.53.00.02.22.02.32.30.610.20.00.10.40.70.40.30.90.02.41.50.12.412.712.8PlotCoder0.00.00.00.00.00.00.06.44.95.00.00.00.03.90.00.00.00.00.00.00.00.00.00.00.00.00.01.61.82.43.02.91.42.62.51.40.40.03.11.80.02.93.06.70.03.56.17.913.40.00.00.011.40.00.00.00.00.00.00.00.0ChDialog1.83.00.81.21.80.52.60.00.00.00.22.00.11.53.00.80.20.10.30.00.00.55.60.72.92.83.90.30.20.00.30.00.00.00.00.00.80.20.15.30.04.92.05.90.01.215.30.01.80.60.91.10.31.30.13.35.60.00.75.65.9Graphics0.02.90.00.00.00.00.00.21.71.83.54.30.69.51.60.80.00.00.00.10.51.05.70.10.20.10.20.00.00.00.00.00.00.00.00.00.01.20.50.20.40.82.35.50.013.921.50.10.31.61.40.81.90.40.61.36.40.50.31.81.5ChartJS2.00.90.60.00.00.014.70.00.10.01.80.00.21.01.50.51.10.10.40.40.40.63.40.73.20.82.00.00.00.00.00.10.00.00.00.01.50.40.10.40.00.71.63.34.61.522.80.20.50.50.50.60.81.30.32.56.00.20.56.46.3Vega-Lite0.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.00.90.03.10.00.00.00.04.40.00.00.00.00.00.00.00.00.00.00.00.045.845.8nvBench
Figure 8: Heat map of the most frequent aesthetic attributes over 7 datasets. The attributes are classified by different categories with colours, such as x and y axes, data appearance, annotation, title and subtitle, legend, grid, figure format, and others.
multiple-plotsmargin-paddingbackgroundbounding-box-bordersaving-formatdpisizegrid-linewidthgrid-linestylegrid-colorgrid-visiblelegend-is-displaylegend-labels-colorlegend-labelslegend-positionlegend-fontsizelegend-titlesubtitle-fontsizesubtitletitle-positiontitle-colortitle-fontsizetitleann-fontann-possitionann-fontsizeann-text-labelerrbar-visibleerrbar-colorerrbar-cap-thickerrbar-cap-sizepie-radiuspie-precision-digitspie-percentage-distancepie-label-distancepie-explodehist-binsbar-data-stackingbar-thicknessmarkerstyleline-capstylelinestylemarkersizelinewidthopacityedge-colorfilled-colorinvert-x-y-axisx-y-scale-positiony-axis-ticks-visiblex-axis-ticks-visiblex-y-ticks-fontsizex-y-scalex-y-ticks-labels-rotationx-y-ticks-labels-colorx-y-ticks-labelsx-y-limx-y-title-colorx-y-title-fontsizey-titlex-titleAttribute70.26.70.50.11.15.550.10.51.00.74.579.50.036.014.93.40.71.63.70.30.36.744.213.263.651.133.6100.011.37.424.57.874.26.34.036.372.85.140.27.60.06.39.510.88.72.333.60.00.90.00.03.32.35.60.412.47.20.87.846.046.3Matplotlib-nb22.98.60.30.11.26.424.20.50.80.83.140.30.016.82.12.20.51.23.70.20.13.326.614.076.747.426.138.115.37.223.59.465.17.55.129.777.99.555.06.6100.07.611.714.59.12.739.30.71.58.612.02.71.82.50.48.76.90.54.730.029.9Matplotlib-py8.40.60.00.00.10.48.70.00.10.10.98.80.016.92.50.20.10.20.50.00.01.212.56.943.836.434.9100.06.90.96.05.477.11.43.144.662.11.931.25.30.05.47.36.76.21.523.50.00.00.00.00.30.41.10.01.90.90.11.515.015.0PlotCoder0.00.00.00.00.00.00.058.343.243.30.00.00.053.20.00.00.00.00.00.00.00.00.00.00.00.00.08.49.712.815.947.422.643.141.123.244.00.036.416.10.016.626.420.40.013.229.538.221.20.00.00.027.60.00.00.00.00.00.00.00.0ChDialog
Figure 9: Heat map of attributes that the user often specifies values when permitted.
Category Attribute Matplotlib R ChartJS Vega-Lite ChartDialog
x-y-axis lim
add_subplot|xlim
set|xlim
axes|xlim
set_xlim|left
set_xlim|right
xlim|left
xlim|right
xlim|xmin
xlim|xmax
set_xlim|xmin
set_xlim|xmax
add_subplot|ylim
set|ylim
axes|ylim
set_ylim|bottom
set_ylim|top
ylim|bottom
ylim|top
ylim|ymin
ylim|ymax
set_ylim|ymin
set_ylim|ymax
plot|xlim
hist|xlim
barplot|xlim
plot.window|xlim
plot.default|xlim
matplot|xlim
boxplot|xlim
curve|xlim
points|xlim
lines|xlim
plot|ylim
barplot|ylim
boxplot|ylim
hist|ylim
lines|ylim
matplot|ylim
points|ylim
plot.window|ylim
plot.default|ylim
options|scales|xAxes|ticks|beginAtZero
options|scales|xAxes|ticks|suggestedMax
options|scales|xAxes|ticks|min
options|scales|xAxes|ticks|max
options|scales|x|beginAtZero
options|scales|x|min
options|scales|x|max
options|scales|yAxes|ticks|beginAtZero
options|scales|y|beginAtZero
options|scales|yAxes|ticks|suggestedMax
options|scales|yAxes|ticks|min
options|scales|yAxes|ticks|max
encoding|x|scale|domain
encoding|x|scale|domain|selection
encoding|x|scale|domain|param
encoding|y|scale|domain
encoding|y|scale|domain|selection
encoding|y|scale|domain|param
Not found
x-y-axis x-y-scale
xscale|value
set_xscale|value
yscale|value
plot|log
lines|log
boxplot|log
options|scales|xAxes|type
options|scales|x|type
options|scales|yAxes|type
options|scales|y|type
encoding|x|scale|type
encoding|y|scale|type
plot|x_axis_scale
contour|x_axis_scale
bar|x_axis_scale
scatter|x_axis_scale
plot|y_axis_scale
contour|y_axis_scale
bar|y_axis_scale
scatter|y_axis_scale
data
appearance
color
plot|color
plot|c
scatter|color
scatter|c
plot|fmt=color
axvline|color
axhline|color
bar|color
barh|color
fill_between|color
hist|color
errorbar|color
contour|colors
set_facecolor|color
vlines|color
hlines|color
Circle|color
Line2D|color
axvspan|color
quiver|color
pie|colors
arrow|color
text|color
annotate|color
lines|col
plot|col
points|col
abline|col
barplot|col
hist|col
polygon|col
barplot|col
rect|col
segments|col
boxplot|col
image|col
curve|col
pie|col
matplot|col
contour|col
stripchart|col
text|col
mtext|col
data|datasets|backgroundColor
data|datasets|pointBackgroundColor
options|plugins|datalabels|color
data|datasets|fillColor
data|datasets|strokeColor
options|elements|line|backgroundColor
data|datasets|pointColor
options|plugins|crosshair|line|color
options|elements|point|backgroundColor
data|datasets|pointStrokeColor
options|plugins|datalabels|color
encoding|color|scale|range
encoding|color|value
encoding|color|aggregate
mark|fill
mark|color
encoding|color|scale|scheme
encoding|color|sort
config|mark|color
mark|color|stops|offset
mark|color|stops|color
encoding|color|type
bar|bar_face_color
hist|bar_face_color
plot|line_color
plot|marker_face_color
scatter|marker_face_color
data
appearance
markerstyle
plot|fmt=markerstyle
plot|marker
scatter|marker
errorbar|marker
points|pch
plot|pch
legend|pch
lines|pch
pairs|pch
matplot|pch
stripchart|pch
par|pch
data|datasets|pointStyle
encoding|shape|type
encoding|shape|value
plot|marker_type
scatter|marker_type
data
appearance
bar-thickness
bar|width
barh|height
barplot|width
data|datasets|barPercentage
data|datasets|maxBarThickness
data|datasets|barThickness
data|datasets|barWidth
options|barThickness
mark|width
mark|height
bar|bar_height
bar|bar_width
data
appearance
hist-bins hist|bins hist|breaks Not found
encoding|x|bin
encoding|x|bin|step
encoding|x|bin|maxbins
encoding|y|bin|maxbins
encoding|x|bin|step
hist|number_of_bins
title title
title|label
set_title|label
set|title
set_title|title
plot|main
hist|main
title|main
barplot|main
boxplot|main
pie|main
image|main
matplot|main
pairs|main
plot.default|main
curve|main
options|title|display
options|title|text
options|plugins|title|display
options|plugins|title|text
options|title
title|title
title|text
Not found
Table 5: Details of the cross-language mapping table for 7 attributes over 3 categories. Each parameter in attributes comprises two parts, function name and argument name, separated by “||||”.